# Markovian Flow Matching: Accelerating MCMC with Continuous Normalizing Flows

### v1

@@ Appendix C Additional Experimental Details
Code for the numerical experiments is written in Python with array computations handled by JAX [9]. The implementation of relevant methods for comparison is sourced from open source repositories: DDS using franciscovargas/denoising_diffusion_samplers, NF-MCMC using kazewong/flowMC, and FAB using lollcat/fab-jax. All experiments are run on an NVidia V100 GPU with 32GB of memory. In the following subsections, we will give more details on the modelling and hyperparameter choices for each experiment, along with additional results.
@@ C.1 Diagnostics
Let $\pi$ and $\nu$ be two probability measures. Let $\mathcal{F}$ denote the unit ball in a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, associated with the positive definite kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. Then the maximum mean discrepancy (MMD) between $\pi$ and $\nu$ is defined as [26, Section 2.2]
@@
where $m_{\pi}$ is the mean embedding of $\pi$, defined via $\mathbb{E}_{\pi}[f]=\langle f,m_{\pi}\rangle_{\mathcal{H}}$ for all $f\in\mathcal{H}$. Using standard properties of the RKHS, the squared MMD can be written as [26, Lemma 6]
@@
Alternatively, we can estimate $\smash{\mathrm{KSD}_{k}^{2}(\pi,\nu)}$ using a biased (but non-negative) V-statistic of the form [43, Section 4]
@@
In all of our numerical experiments, we calculate the U- and V- statistics using the inverse multi-quadratic kernel $k(x,x^{\prime})=(1+(x-x^{\prime})^{T}(x-x^{\prime}))^{\beta}$ due to its favourable convergence properties [23, Theorem 8], setting $\beta=-\frac{1}{2}$.
@@
In Table 3, we present results for $K=10^{3}$ iterations.Since MFM is much more efficient than other methods, we also report results for a great number of total iterations. Table 4 contains results for $K=5\cdot 10^{3}$ iterations for MFM. In the main text, we present results for $K=5\cdot 10^{3}$ learning iterations for MFM, and $K=10^{3}$ iterations for the other algorithms, since this renders the total computational cost of all algorithms somewhat comparable.
@@
Like the 4-mode example, all methods use $N=128$ parallel chains for training and $128$ hidden dimensions for all neural networks. Methods with a MALA kernel use a step size of $0.2$, and methods with splines use 4 coupling layers with 8 bins and range limited to $[-16,16]$. In Table 5 we present results for $K=10^{3}$ iterations. In Table 6, we provide results for MFM for $K=5\times 10^{3}$ learning iterations. In the main text, we present results for $K=5\cdot 10^{3}$ learning iterations for MFM and $K=10^{3}$ iterations for all other algorithms, which yields a more comparable total computation time.
@@
C.4 Field system
The Allen–Cahn equation is defined in terms of a random field $\phi:[0,1]\rightarrow\mathbb{R}$ satisfying the following stochastic partial differential equation:
@@ C.4 Field system
where $a>0$ is a parameter, $\beta$ is the inverse temperature, $s\in[0,1]$ denotes the spatial variable, and $\eta$ is spatiotemporal white noise. The imposition of Dirichlet boundary conditions ensures $\phi(s=0)=\phi(s=1)=0$.
@@
C.5 Log-Gaussian Cox point process
The original $10\times 10$ square meter plot is standardized to the unit square, as depicted in Fig. 4. We discretize the unit square $[0,1]^{2}$ into a $M=40\times 40$ regular grid. The latent intensity process $\Lambda=\{\Lambda_{m}\}_{m\in M}$ is specified as $\Lambda_{m}=\exp(X_{m})$, where $X=\{X_{m}\}_{m\in M}$ is a Gaussian process with a constant mean $\mu_{0}\in\mathbb{R}$ and exponential covariance function $\Sigma_{0}(m,n)=\sigma^{2}\exp\left(-|m-n|/(30\beta)\right)$ for $m,n\in M$, i.e. $X\sim\mathcal{N}(\mu_{0}1_{d},\Sigma_{0})$ for $1_{d}=[1,\dots,1]^{T}\in\mathbb{R}^{d}$ with dimension $d=900$. The chosen parameter values are $\sigma^{2}=1.91$, $\beta=1/33$, and $\mu_{0}=\log(126)-\sigma^{2}/2$, estimated by [50]. The number of points in each grid cell $Y=\{Y_{m}\}_{m\in M}\in\mathbb{N}^{40\times 40}$ are modelled as conditionally independent and Poisson distributed with means $a\Lambda_{m}$,
@@ C.5 Log-Gaussian Cox point process
where $a=1/40^{2}$ represents the area of each grid cell.
Figure 4: Visualization of the standardized 10 $\times$ 10 square meter plot and the locations of 126 Scots pine saplings in Finland.
For this example, all methods use $N=128$ parallel chains for training and $1024$ hidden dimensions for all neural networks. Methods with a MALA kernel use a step size of $0.01$, and methods with splines use 8 coupling layers with 8 bins and range limited to $[-10,10]$. Results for $K=10^{4}$ learning iterations are presented in Table 8. In the main text, we present results for $k_{Q}=10^{3}$ for MFM. We were unable to run NF-MCMC and FAB for $K=10^{4}$ iterations because of memory issues; instead, we present results for $K=10^{3}$ iterations only for the models using discrete normalizing flows.

### v2

@@ Appendix C Additional Experimental Details
Code for the numerical experiments is written in Python with array computations handled by JAX [11]. The implementation of relevant methods for comparison is sourced from open source repositories: DDS using franciscovargas/denoising_diffusion_samplers, NF-MCMC using kazewong/flowMC [80], and FAB using lollcat/fab-jax [52]. All experiments are run on an NVIDIA V100 GPU with 32GB of memory. In the following subsections, we will give more details on the modelling and hyperparameter choices for each experiment, along with additional results.
@@ C.1 Diagnostics
Let $\pi$ and $\nu$ be two probability measures. Let $\mathcal{F}$ denote the unit ball in a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, associated with the positive definite kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. Then the maximum mean discrepancy (MMD) between $\pi$ and $\nu$ is defined as [29, Section 2.2]
@@
where $m_{\pi}$ is the mean embedding of $\pi$, defined via $\mathbb{E}_{\pi}[f]=\langle f,m_{\pi}\rangle_{\mathcal{H}}$ for all $f\in\mathcal{H}$. Using standard properties of the RKHS, the squared MMD can be written as [29, Lemma 6]
@@
Alternatively, we can estimate $\smash{\mathrm{KSD}_{k}^{2}(\pi,\nu)}$ using a biased (but non-negative) V-statistic of the form [46, Section 4]
@@
In all of our numerical experiments, we calculate the U- and V- statistics using the inverse multi-quadratic kernel $k(x,x^{\prime})=(1+(x-x^{\prime})^{T}(x-x^{\prime}))^{\beta}$ due to its favourable convergence properties [26, Theorem 8], setting $\beta=-\frac{1}{2}$.
@@
In Table 3, we present results for $K=10^{3}$ iterations. Since MFM is much more efficient than other methods, we also report results for a great number of total iterations. Table 4 contains results for $K=5\cdot 10^{3}$ iterations for MFM and AT-SMC. In the main text, we present results for $K=5\cdot 10^{3}$ learning iterations for MFM and AT-SMC, and $K=10^{3}$ iterations for the other algorithms, since this renders the total computational cost of all algorithms somewhat comparable.
@@
Like the 4-mode example, all methods use $N=128$ parallel chains for training and $128$ hidden dimensions for all neural networks. Methods with a MALA kernel use a step size of $0.2$, and methods with splines use 4 coupling layers with 8 bins and range limited to $[-16,16]$. In Table 5 we present results for $K=10^{3}$ iterations. In Table 6, we provide results for MFM and AT-SMC for $K=5\times 10^{3}$ learning iterations. In the main text, we present results for $K=5\cdot 10^{3}$ learning iterations for MFM and AT-SMC and $K=10^{3}$ iterations for all other algorithms, which yields a more comparable total computation time.
@@
C.4 Many Well
We also present a synthetic problem approximating the 32-dimensional Many Well distribution given by the product of 16 copies of the 2-dimensional Double Well distribution [e.g., 81, 58, 52],
$\log p(x_{1},x_{2})=-x_{1}^{4}+6x_{1}^{2}+\frac{1}{2}x_{1}-\frac{1}{2}x_{2}^{2}+\mathrm{constant},$ (41)
where each copy of the Double Well is evaluated on a different pair of the 32 inputs. The 32-dimensional Many Well has $655362^{16}=$ 65536 modes, one for each possible choice of mode in each of the 16 copies of the double well. We can obtain exact samples from the Many Well by sampling each independent copy of the Double Well.
Like previous synthetic examples, all methods use $N=128$ parallel chains for training and $128$ hidden dimensions for all neural networks. Methods with a MALA kernel use a step size of $0.1$, and methods with splines use 4 coupling layers with 8 bins and a range limited to $[-16,16]$. Table 7 presents results for $K=10^{3}$ iterations. In Table 8, we provide results for MFM and AT-SMC for $K=5\times 10^{3}$ learning iterations. In the main text, we present results for $K=5\cdot 10^{3}$ learning iterations for MFM and AT-SMC and $K=10^{3}$ iterations for all other algorithms, which yields a more comparable total computation time.
FM w/ $\pi$ samples $86.6\pm 1.88$ $17.4\pm 6.24$ $19.3\pm 6.25$ $1.24$e-8±$\pm 1.81$e-8 $6.95\pm 0.31$
MFM $k_{Q}=K$ $101.\pm 0.70$ $0.12\pm 0.12$ $0.77\pm 0.13$ $2.28$e-8±$\pm 1.57$e-8 $11.1\pm 0.67$
MFM $k_{Q}=10^{2}$ $101.\pm 0.70$ $0.12\pm 0.11$ $0.77\pm 0.13$ $2.28$e-8±$\pm 1.57$e-8 $120.\pm 17.6$
– w/ HTE $101.\pm 0.70$ $0.11\pm 0.12$ $0.77\pm 0.13$ $2.28$e-8±$\pm 1.57$e-8 $35.6\pm 3.95$
MFM $k_{Q}=10$ $101.\pm 0.77$ $0.11\pm 0.11$ $0.76\pm 0.12$ $2.28$e-8±$\pm 1.57$e-8 $1100.\pm 90.2$
– w/ HTE $101.\pm 0.69$ $0.11\pm 0.09$ $0.76\pm 0.10$ $2.27$e-8±$\pm 1.56$e-8 $259.\pm 17.6$
DDS $133.\pm 3.92$ $0.13\pm 0.14$ $0.61\pm 0.13$ $1.49$e-5±$\pm 2.03$e-5 $227.\pm 0.91$
NF-MCMC $39.2\pm 1.61$ $1.17\pm 0.65$ $2.29\pm 0.66$ $4.79$e-7±$\pm 2.41$e-7 $184.\pm 44.1$
FAB $137.\pm 0.33$ $4.79$e-3±$\pm 2.55$e-2 $4.32$e-1±$\pm 4.14$e-2 $3.87$e-7±$\pm 2.90$e-7 $304.\pm 3.64$
AT-SMC $130.\pm 2.93$ $0.02\pm 0.03$ $0.57\pm 0.03$ $1.75$e-5±$\pm 1.19$e-5 $1.35\pm 0.36$
Table 7: Diagnostics for the many well with $K=10^{3}$. $\mathbb{E}_{[\phi_{1}]_{\#}p_{0}}\log\pi$ is the Monte Carlo approximation of the log-target density using the learned flow to generate samples; KSD U-stat and V-stat are the Kernel Stein Discrepancy U- and V-statistics between the target and samples generated from the learned flow; MMD is the Maximum Mean Discrepancy between real samples from the target and samples generated from the learned flow. Results are averaged and empirical 95% confidence intervals over 10 independent runs.
FM w/ $\pi$ samples $98.6\pm 5.32$ $2.57\pm 4.07$ $4.56\pm 4.07$ $-1.13$e-9±$\pm 2.81$e-8 $25.6\pm 0.48$
MFM $k_{Q}=K$ $101.\pm 0.64$ $1.02$e-1±$\pm 9.03$e-2 $7.64$e-1±$\pm 9.78$e-2 $2.28$e-8±$\pm 1.57$e-8 $29.7\pm 1.03$
MFM $k_{Q}=10^{2}$ $101.\pm 0.63$ $1.02$e-1±$\pm 9.05$e-2 $7.64$e-1±$\pm 9.73$e-2 $2.28$e-8±$\pm 1.57$e-8 $587.\pm 27.0$
– w/ HTE $102.\pm 0.78$ $1.02$e-1±$\pm 8.32$e-2 $7.62$e-1±$\pm 8.91$e-2 $2.27$e-8±$\pm 1.58$e-8 $154.\pm 7.51$
MFM $k_{Q}=10$ $102.\pm 0.81$ $9.93$e-2±$\pm 7.24$e-2 $7.63$e-1±$\pm 8.15$e-2 $2.27$e-8±$\pm 1.57$e-8 $5130.\pm 104$
– w/ HTE $103.\pm 2.20$ $3.85$e-1±$\pm 2.69$e-1 $1.05\pm 0.27$ $2.21$e-8±$\pm 1.57$e-8 $1190.\pm 27.6$
AT-SMC $130.\pm 3.17$ $2.44$e-2±$\pm 5.99$e-2 $5.79$e-1±$\pm 6.76$e-2 $1.52$e-5±$\pm 1.00$e-5 $2.59\pm 0.27$
Table 8: Diagnostics for the many well with $K=5\cdot 10^{3}$. $\mathbb{E}_{[\phi_{1}]_{\#}p_{0}}\log\pi$ is the Monte Carlo approximation of the log-target density using the learned flow to generate samples; KSD U-stat and V-stat are the Kernel Stein Discrepancy U- and V-statistics between the target and samples generated from the learned flow; MMD is the Maximum Mean Discrepancy between real samples from the target and samples generated from the learned flow. Results are averaged and empirical 95% confidence intervals over 10 independent runs.
C.5 Field system
The stochastic Allen–Cahn equation is defined in terms of a random field $\phi:[0,1]\rightarrow\mathbb{R}$ satisfying the following stochastic partial differential equation [e.g., 24, Section V]:
@@ C.5 Field system
where $a>0$ is a parameter, $\beta$ is the inverse temperature, $s\in[0,1]$ denotes the spatial variable, and $\eta$ is spatiotemporal white noise. We impose Dirichlet boundary conditions throughout, so that $\phi(s=0)=\phi(s=1)=0$.
@@
C.6 Log-Gaussian Cox process
The original $10\times 10$ square meter plot is standardized to the unit square. We discretize the unit square $[0,1]^{2}$ into a $M=40\times 40$ regular grid. The latent intensity process $\Lambda=\{\Lambda_{m}\}_{m\in M}$ is specified as $\Lambda_{m}=\exp(X_{m})$, where $X=\{X_{m}\}_{m\in M}$ is a Gaussian process with a constant mean $\mu_{0}\in\mathbb{R}$ and exponential covariance function $\Sigma_{0}(m,n)=\sigma^{2}\exp\left(-|m-n|/(40\beta)\right)$ for $m,n\in M$, i.e. $X\sim\mathcal{N}(\mu_{0}1_{d},\Sigma_{0})$ for $1_{d}=[1,\dots,1]^{T}\in\mathbb{R}^{d}$ with dimension $d=1600$. The chosen parameter values are $\sigma^{2}=1.91$, $\beta=1/33$, and $\mu_{0}=\log(126)-\sigma^{2}/2$, corresponding to the values estimated in [53]. The number of points in each grid cell $Y=\{Y_{m}\}_{m\in M}\in\mathbb{N}^{40\times 40}$ are modelled as conditionally independent and Poisson distributed with means $a\Lambda_{m}$,
@@ C.6 Log-Gaussian Cox process
where $a=1/40^{2}$ represents the area of each grid cell. For this example, all methods use $N=128$ parallel chains for training and $1024$ hidden dimensions for all neural networks. Methods with a MALA kernel use a step size of $0.01$, and methods with splines use 8 coupling layers with 8 bins and range limited to $[-10,10]$. Results for $\smash{K=10^{4}}$ learning iterations are presented in Table 10. In the main text, we present results for $\smash{k_{Q}=10^{3}}$ for MFM. We were unable to run NF-MCMC and FAB for $\smash{K=10^{4}}$ iterations because of memory issues; instead, we present results for $\smash{K=10^{3}}$ iterations only for the models using discrete normalizing flows.